### v1

@@ 4.1 NARMA: imitating known dynamics
NARMA (Normalised Auto-Regressive Moving Average) is a family of imitation tasks widely used as benchmarks. It is based on two systems for modelling time series based dynamical systems (Weigend and Gershenfeld,, 1994), the Auto-Regressive (AR) and Moving Average (MA) models.
Connor et al., (1991, 1994) introduce and define a generalisation of linear ARMA models; their generic NARMA$(p,q)$ model depends on $p$ past states and $q$ past inputs:
$x(t)$ $=h(x(t-1),\ldots,x(t-p),$ (3)
$\qquad u(t-1),\ldots,u(t-q))+u(t)$
@@
Atiya and Parlos, (2000, eqn.79) introduce a second order NARMA system:
$x(t+1)$ $=0.4x(t)+{}$ (4)
$0.1\displaystyle\quad 0.4x(t)x(t-1)+0.6u^{3}(t)+$ 0.1
This is sometimes referred to in the literature as ‘NARMA2’ (Fujii and Nakajima,, 2017), but is not of the same functional form as the family of equations usually referred to as NARMA-$N$.
Atiya and Parlos, (2000, eqn.86) also introduce a tenth order NARMA system, which is now referred to as NARMA-10. The functional form has been generalised (Rodan and Tino,, 2011) as the $N$th order, or NARMA-$N$, system:
$x(t+1)$ $=\alpha x(t)+{}$ (5)
$\qquad\beta x(t)\left(\sum_{i=0}^{N-1}x(t-i)\right)+{}$
$\qquad\gamma u(t-N+1)u(t)+\delta$
@@
NARMA-10 using the original parameter values and the input stream generated from a uniform random distribution in the range [0, 0.5] is one of the commonest usages in the literature; see for example Verstraeten et al., (2007), Rodan and Tiňo, (2010), Holzmann and Hauser, (2010), Paquot et al., (2012), Goudarzi et al., (2015) Vinckier et al., (2015), Inubushi and Yoshimura, (2017), Dale, 2018a .
@@
5 0.3 0.05 1.5 0.1 $U[0,0.2]$ Fujii and Nakajima, (2017, eqn.18)
5 0.3 0.05 1.5 0.1 $U[0,0.5]$ Dale, 2018b (, chap.3)
10 0.3 0.05 1.5 0.1 $U[0,0.5]$ Atiya and Parlos, (2000, eqn.86)
15 0.3 0.05 1.5 0.1 $U[0,0.2]$ Fujii and Nakajima, (2017, eqn.18)
20 0.3 0.05 1.5 0.01 $U[0,0.5]$ Rodan and Tino, (2011, eqn.6), with tanh(.)
20 0.3 0.05 1.5 0.1 $U[0,0.2]$ Fujii and Nakajima, (2017, eqn.18)
30 0.2 0.04 1.5 0.001 $U[0,0.5]$ Schrauwen et al., 2008b (, sec.3)
30 0.2 0.004 1.5 0.001 $U[0,0.5]$ Dale, 2018a
Table 2: Summary of NARMA-$N$ parameters found in literature. The NARMA sequence is plotted for the starred values in Fig.4
Adapting this for different values of $N$ reveals that the equation can rapidly diverge if not tuned correctly. Indeed, even the standard NARMA-10 setup does itself occasionally diverge (Kubota et al.,, 2021). Other authors have defined the form for further values of $N$, and taken a variety of approaches to controlling divergence. For example, Schrauwen et al., 2008b (, p.1164) defines NARMA-30 with different parameter values of $(0.2,0.04,1.5,0.001)$, nevertheless, it too very occasionally diverges); Rodan and Tino, (2011, eqn.6) define NARMA-20 and add a $\tanh(.)$ wrapper to stop divergence; Fujii and Nakajima, (2017, eqn.18) define NARMA-5, 10, 15, and 20 with the original parameter values but a restricted input range $u\in\mbox{{Uniform}}[0,0.2]$. Parameter values for these and other cases from the literature are given in table 2.
Furthermore, various NARMA time series each have their own range of output $x$ values. For example, Atiya and Parlos, (2000)’s NARMA-10 $x$ values range between (ignoring divergences) 0.15 and 1, Fujii and Nakajima, (2017)’s NARMA-10 with reduced $u$ range has $x$ values ranging between 0.15 and 0.25, and Schrauwen et al., 2008b ’s NARMA-30 $x$ values range between (again ignoring divergences) 0 and 0.6. Even using consistent parameter values and input range yields systematically different output ranges for different values of $N$. This makes any performance comparison as a function of $N$ problematic.
The relative simplicity of the task, and the ability to parameterise the difficulty by using a range of $N$ values, has made this benchmark popular when evaluating small reservoirs, as is often true for in materio reservoir computers (Dale, 2018b, ). As noted, however, however, there are several issues: the potential for divergence when generating the series, a multitude of parameter values in the literature making comparisons difficult, and no clear trends as a function of $N$. As such, its use as a benchmark should be approached with caution. NARMA is a non-linear memory task, so using the generic, more readily interpretable, albeit more computationally intensive, Information Processing Capacity (IPC) measure (section 8.1.2) can overcome these limitations.
Authors using this benchmark typically use it alongside other benchmarks, such as the Santa Fe Laser Task (Inubushi and Yoshimura,, 2017) (sec.5.7), the Mackey–Glass System (Goudarzi et al.,, 2015) (sec.5.1) or the Multiple Superimposed Sines task (Holzmann and Hauser,, 2010) (sec.5.2).
@@ 4.2 Channel equalisation: imitating unknown dynamics
Channel Equalisation can be modelled as an imitation task: the Reservoir Computer imitates a “perfect” filter, which completely removes the noise. This filter is unknown and may even be impossible, but as the output is known (the noiseless signal) and the training relies solely on the output, it can still be imitated to some degree. This task may be evaluated using a Symbol Error Rate (Jaeger and Haas,, 2004).
Jaeger and Haas, (2004) first applied the task to Reservoir Computing. A description of the method used can be found in their supplementary materials (Jaeger, 2004a, ). The channel model used by Jaeger and Haas, (2004) was first introduced by Mathews and Lee, (1994).
The system takes as input a signal to which nonlinear noise has been added: the task is to output the signal with the noise removed. Jaeger, 2004a use a signal $d(n)$ that is a sequence of randomly chosen values from the set $\{-3,-1,1,3\}$. The signal $d(n)$ is passed through a linear channel [^2] In Jaeger, 2004a the coefficient is written as 0.09 1; however, both the original (Mathews and Lee,, 1994) and the “other formulation” (Rodan and Tiňo,, 2010) use 0.09, indicating that this may be a typographical error.:
$q(n)$ $=0.08d(n+2)-0.12d(n+1)+{}$ (6)
$d(n)+0.18d(n-1)-0.1d(n-2)+{}$
$0.09d(n-3)-0.05d(n-4)+0.04d(n-5)+{}$
$0.03d(n-6)+0.01d(n-7)$
@@
where $v(n)$ is Gaussian white noise, applied to ensure that the signal-to-noise ratio of the output is between 12 and 32 decibels (Jaeger and Haas,, 2004). [^3] Jaeger and Haas, (2004) specify 12 to 32 dB; the supplementary material (Jaeger, 2004a, ) specifies 16 to 32 dB. Most authors who use the benchmark (Vinckier et al.,, 2015; Paquot et al.,, 2012; Dejonkheere et al.,, 2014; Duport et al., 2012a, ) use the lower bound of 12 dB. There are some other formulations; for example Rodan and Tiňo, (2010) have no noise term, and add an offset of +30. An example of such a signal, as well as the signal with noisy applied, can be found in figure 5.
Figure 5: A channel signal (blue), and the signal with nonlinear noise applied (orange).
One can also find certain tasks in the literature that are called “Channel Equalisation”, but do not follow the same methodology as the best known version of the benchmark. For example, Boccato et al., (2011) uses a similar task, but the signal is composed solely of the values –1, +1, and the evaluation is done via Average Mean Squared Error. Boccato et al., (2012) follows this up by adding an unsupervised channel equalisation task, where the noiseless signal is not known. While this task has a similar purpose to the classic, supervised version of the task, the lack of “clean” signal used in training signifies that it is not an imitation task. This provides us with an interesting distinction between the form of the task (imitation of an unknown dynamical system, or not) and its function (channel equalisation, in both cases).
@@ 4.2.1 Unsupervised Channel Equalisation
An unsupervised version of the Channel Equalisation task exists (Boccato et al.,, 2012). The argument is that most distorted signals do not have a “clean” version that can be used to train the nonlinear filter, therefore it is useful to have a version of the task that uses unsupervised learning as opposed to imitation-based learning. Instead of training the received signal against the perfect, noiseless signal, the reservoir was trained against the information conveyed. The authors argue that, since the explicit signal itself was not used in training, this qualifies the task as an unsupervised learning task. As the ‘perfect signal’ is still known, however, the task is evaluated against that perfect signal.
@@ 4.3 Two-jointed arm: imitating unknown dynamics
The two-jointed arm task (Slotline and Li,, 1991) involves controlling a two-jointed (robotic) arm in order to move it from point A to point B. Like the Channel Equalisation task, this task can be seen as an imitation of an unknown, perfect controller, of which only the outputs are known.
Joshi and Maass, (2004, 2005) use this task in the context of Reservoir Computing. The training data used are points along the trajectory that the robotic arm was intended to follow. The robotic arm is then tested in a “closed loop”, where no training data is available, what we in section 3.1.1 refer to as “in-situ” testing.
@@
where $\boldsymbol{\theta}=[\theta_{1},\theta_{2}]^{T}$ are the angles of the arm’s joints, $\boldsymbol{\tau}=[\tau_{1},\tau_{2}]^{T}$ are the joint input torques, $\mathbf{H}$ is a $2\times 2$ inertia matrix, and $\mathbf{C}$ is a $2\times 2$ matrix of the Coriolis and centripetal terms. The reservoir is then given a target trajectory, for which it must imitate a “perfect” controller in a model-free manner. See Joshi and Maass, (2004, 2005) for details.
@@ 4.4 Van der Pol oscillator: imitating unknown dynamics of a controller
The Van der Pol (VdP) oscillator was first described by the Dutch physicist Balthazar Van der Pol in the 1920s (van der Pol,, 1922) and later employed to model the oscillations of the heart (Van Der Pol and Van Der Mark,, 1928). It is particularly useful for modelling systems that exhibit non-linear behaviour, such as relaxation oscillations and self-sustained oscillations. Within the context of machine learning, the VdP oscillator is commonly used to test the effectiveness of neural networks (Nicola and Clopath,, 2017; Yeo,, 2019). Shougat and Perkins, (2023) apply the VdP oscillator as the substrate of physical reservoir computing.
@@
where $x$ is the state variable, $t$ is time, and $\mu$ is a scalar parameter that dictates the nonlinearity and the strength of the damping. $F(t)$ is the external force applied to the VdP system. In the vanilla VdP system the external force term is zero $F(t)=0$, in which case the phase plane trajectory of the VdP is as shown in Figure 6).
@@
In a recent study, the VdP oscillator is used as a model-free control task where a reservoir is trained to produce an external force to make the VdP system produce a circle (Gan,, 2023; Gan et al.,, 2024).
@@ 4.5 Pole balancing : unsupervised learning
In the field of Reservoir Computing, this task is primarily used to evaluate unsupervised learning in reservoirs. It has been used for exploring different methods of training reservoirs, such as Neuroevolution (Jiang et al., 2008b, ; Jiang et al., 2008a, ) and NEAT (Chatzidimitriou and Mitkas,, 2010). One could also implement this task as an imitation task, and have the reservoir imitate the unknown dynamics of a perfect controller.
@@ 5 Dynamics prediction tasks
In this section, we describe the systems commonly used as prediction benchmarks in Reservoir Computing. A greater variety of datasets that can be used for benchmark tasks can be found in Gilpin, (2021)’s database of chaotic dynamical systems, all specified according to guidelines proposed by Gebru et al., (2021).
@@ 5.1 Mackey–Glass equation : known dynamics
The Mackey–Glass benchmark is derived from a delay differential equation introduced to model certain physiological systems (Mackey and Glass,, 1977). [^5] This delay differential equation is also used elsewhere in Reservoir Computing to create a nonlinear node in delay line reservoirs (see, for example, Appeltant et al., (2011)). This other usage is not examined here. These models characterise illness through changes to the variables that lead to these systems turning chaotic. The authors examine various models, one for the breathing patterns of patients with Cheyne-Stokes illnesses, and two variants for the growth rates of white blood cells in leukaemia patients.
The latter model is used as the basis for the benchmark; Mackey and Glass, (1977, eqn.4b) state:
@@
Normalising the concentration with respect to the baseline, $x_{t}=P_{t}/\theta$, gives the more usual form of the Mackey–Glass chaotic equation (Glass and Mackey,, 2010, eqn.1):
@@
Here $x$ is a dimensionless normalised quantity, expressed in ‘units’ of the parameter $\theta$. $\theta$ is still a parameter of the model, but is now implicit in the equation: it provides the scale for the dependent variable $x$. Eqn.11 has chaotic dynamics for $\beta=2$, $n=9.65$, $\gamma=1$, $\tau=2$ with an initial condition of $x=0.5$ for $t<0$ (Glass and Mackey,, 2010). [^6] The text of (Glass and Mackey,, 2010) states $\beta=0.2,\gamma=0.1,\tau=2$, but the captions of its figures 1 and 2 state $\beta=2,\gamma=1,\tau=2$. Figure 7 here favours the values from the captions. Equivalently, one could use $\beta=0.2,\gamma=0.1,\tau=20$.
@@
Top: $[\beta,n,\gamma,\tau]$ values from Glass and Mackey, (2010):
$[0.2,9.65,0.1,2]$ : transient to a point attractor
$[2,8.5,1,2]$ : periodic behaviour [their fig.5]
$[2,9.65,1,2]$ : chaotic behaviour [their fig.2,fig.7]
Bottom: $[\beta,n,\gamma,\tau]$ values from Jaeger, (2010):
$[0.2,10,0.1,16]$ : periodic behaviour
$[0.2,10,0.1,17]$ : mildly chaotic behaviour
$[0.2,10,0.1,30]$ : wildly chaotic behaviour
@@
Jaeger, (2010) states that the Mackey–Glass system is often used for learning dynamical systems from data, ‘invariably’ with parameter values of $\beta=0.2,n=10,\gamma=0.1$, and that with these values the system is chaotic for $\tau>16.8$; see (Farmer,, 1982) for a detailed analysis. Jaeger, (2010) investigates $\tau=17$ (mildly chaotic) and $\tau=30$ (which he dubs ‘wildly’ chaotic). See figure 7.
The parameter values used by Glass and Mackey, (2010) and by Jaeger, (2010) appear to differ considerably (figure 7); however, timescale rescaling shows they are comparable. There are three timescales in eqn.11: $1/\beta$ (production, or growth, timescale); $1/\gamma$ (decay timescale), and $\tau$ (time delay). If we normalise these with respect to the decay timescale, we have two independent timescales: $\gamma/\beta$ and $\gamma\tau$. Glass and Mackey, (2010) use $\gamma/\beta=1/2$, $\gamma\tau=2$; Jaeger, (2010) uses $\gamma/\beta=1/2$, $\gamma\tau=1.7,3.0$.
Glass and Mackey, (2010) explore the transition between periodic and chaotic behaviour using fixed timescales and varying the non-linearity parameter $n$ (the timescales have biological meaning whereas the non-linearity is a parameter of the model); Jaeger, (2010) fixes the non-linearity at $n=10$ and instead varies the delay feedback timescale $\tau$ (the model is being used in a more abstract manner, and varying the time delay is a standard approach to investigating chaotic behaviours in DDEs (Otto et al.,, 2019)).
Eqn.11 describes a continuous system. Jaeger, (2010) describes a discretisation process for converting it to a discrete system suitable for generating time series data. Discretise eqn.11 using $dx/dt\approx(x(t+\Delta t)-x(t))/\Delta t$:
@@
Take $1/\Delta t=N$, an integer (Jaeger, (2010) uses $N=10$), to produce the time series $\hat{x}(t),\hat{x}(t+\Delta t),\hat{x}(t+2\Delta t),\ldots\hat{x}(t+(N-1)\Delta t),\hat{x}(t+1),\ldots$
@@
Using this discretisation method, Jaeger, (2010) generates two sequences for each value of $\tau=17$ and $30$, one sequence of length 3,000 and of 21,000 (four sequences in total). These form the combined washout and training sequences. The ESN is trained on this data, then tested on its ability to predict the 84th output value, compared to the expected 84th value, computing the relevant Normalised Mean Squared Error (NMSRE84) of the value over multiple runs.
Mackey–Glass is used as a benchmark by Jaeger again in (Jaeger and Haas,, 2004), this time with a single 3,000-step training sequence, and the NMSRE of the 84th subsequent value taken over 100 runs. The $\tau$ values used are not specified. Holzmann and Hauser, (2010) also use the system as a benchmark, using a method almost identical to (Jaeger,, 2010), though with the NMSRE taken over 100 runs, and an additional NMSRE of the 120th value used for the training sequence of length 21,000. Roeschies and Igel, 2010a base their method on (Jaeger,, 2010), this time using a single sequence of length 3,000 with $\tau=17$. One other notable difference is the use of a washout of 100 values, whereas prior works have used washouts of 1,000. Goudarzi et al., (2015) and Moon et al., (2019) depart from this traditional setup in order to propose their own. These various different uses are detailed in table 3.
@@
Roeschies and Igel, 2010a 1 3000 17 100 84 50
Goudarzi et al., (2015) 1 2000 ? ? 2000 10
Moon et al., (2019) 1 500 18 ? 50–52 ?
[…]
Table 3: A summary of different parameters used in the literature for the Mackey–Glass prediction benchmark. Values marked with “?” are not reported in the cited work. Goudarzi et al., (2015) use NMSE rather than NMSRE.

### v2

@@ 4.1 NARMA: imitating known dynamics
NARMA (Nonlinear Auto-Regressive Moving Average) is a family of imitation tasks widely used as benchmarks. It is based on two systems for modelling time series based dynamical systems [32], the Auto-Regressive (AR) and Moving Average (MA) models.
Connor et al. [33, 34] introduce and define a generalisation of linear ARMA models; their generic NARMA$(p,q)$ model depends on $p$ past states and $q$ past inputs:
$x(t)$ $=h(x(t-1),\ldots,x(t-p),u(t-1),\ldots,u(t-q))+u(t)$ (3)
@@
Atiya and Parlos [35, eqn.79] introduce a second order NARMA system:
$x(t+1)$ $=0.4x(t)+{}0.4x(t)x(t-1)+0.6u^{3}(t)+0.1$ (4)
This is sometimes referred to in the literature as ‘NARMA2’ [36], but is not of the same functional form as the family of equations usually referred to as NARMA-$N$.
Atiya and Parlos [35, eqn.86] also introduce a tenth order NARMA system, which is now referred to as NARMA-10. The functional form has been generalised [37] as the $N$th order, or NARMA-$N$, system:
$x(t+1)$ $=\alpha x(t)+{}\beta x(t)\left(\sum_{i=0}^{N-1}x(t-i)\right)+{}\gamma u(t-N+1)u(t)+\delta$ (5)
@@
NARMA-10 using the original parameter values and the input stream generated from a uniform random distribution in the range [0, 0.5] is one of the commonest usages in the literature; see for example [38, 14, 39, 40, 41, 42, 43, 44].
@@
5 0.3 0.05 1.5 0.1 $U[0,0.2]$ [36, eqn.18]
5 0.3 0.05 1.5 0.1 $U[0,0.5]$ [45, chap.3]
10 0.3 0.05 1.5 0.1 $U[0,0.5]$ [35, eqn.86]
15 0.3 0.05 1.5 0.1 $U[0,0.2]$ [36, eqn.18]
20 0.3 0.05 1.5 0.01 $U[0,0.5]$ [37, eqn.6], with tanh(.)
20 0.3 0.05 1.5 0.1 $U[0,0.2]$ [36, eqn.18]
30 0.2 0.04 1.5 0.001 $U[0,0.5]$ [46, sec.3]
30 0.2 0.004 1.5 0.001 $U[0,0.5]$ [44]
Table 2: Summary of NARMA-$N$ parameters found in literature. The NARMA sequence is plotted for the starred values in Figure 4
Adapting this for different values of $N$ reveals that the equation can rapidly diverge if not tuned correctly. Indeed, even the standard NARMA-10 setup does itself occasionally diverge [47]. Other authors have defined the form for further values of $N$, and taken a variety of approaches to controlling divergence. For example, Schrauwen et al. [46, p.1164] define NARMA-30 with different parameter values of $(0.2,0.04,1.5,0.001)$, nevertheless, it too very occasionally diverges); Rodan and Tino [37, eqn.6] define NARMA-20 and add a $\tanh(.)$ wrapper to stop divergence; Fujii and Nakajima [36, eqn.18] define NARMA-5, 10, 15, and 20 with the original parameter values but a restricted input range $u\in\mbox{{Uniform}}[0,0.2]$. Parameter values for these and other cases from the literature are given in table 2.
Furthermore, various NARMA time series each have their own range of output $x$ values. For example, Atiya and Parlos [35]’s NARMA-10 $x$ values range between (ignoring divergences) 0.15 and 1, Fujii and Nakajima [36]’s NARMA-10 with reduced $u$ range has $x$ values ranging between 0.15 and 0.25, and d Schrauwen et al. [46]’s NARMA-30 $x$ values range between (again ignoring divergences) 0 and 0.6. Even using consistent parameter values and input range yields systematically different output ranges for different values of $N$. This makes any performance comparison as a function of $N$ problematic.
The relative simplicity of the task, and the ability to parameterise the difficulty by using a range of $N$ values, has made this benchmark popular when evaluating small reservoirs, as is often true for in materio reservoir computers [45]. As noted, however, there are several issues: the potential for divergence when generating the series, a multitude of parameter values in the literature making comparisons difficult, and no clear trends as a function of $N$. As such, its use as a benchmark should be approached with caution. NARMA is a non-linear memory task, so using the generic, more readily interpretable, albeit more computationally intensive, Information Processing Capacity (IPC) measure (section 8.1.2) can overcome these limitations.
Authors using this benchmark typically use it alongside other benchmarks, such as the Santa Fe Laser Task [43] (sec.5.7), the Mackey–Glass System [41] (sec.5.1) or the Multiple Superimposed Sines task [39] (sec.5.2).
@@ 4.2 Channel equalisation: imitating unknown dynamics
Channel Equalisation can be modelled as an imitation task: the Reservoir Computer imitates a ‘perfect’ filter, which completely removes the noise. This filter is unknown and may even be impossible, but as the output is known (the noiseless signal) and the training relies solely on the output, it can still be imitated to some degree. This task may be evaluated using the symbol error rate [25].
Jaeger and Haas [25] first applied the task to Reservoir Computing. A description of the method used can be found in their supplementary materials [48]. The channel model used in [25] was first introduced by Mathews and Lee [49].
The system takes as input a signal to which nonlinear noise has been added: the task is to output the signal with the noise removed. Jaeger [48] uses a signal $d(n)$ that is a sequence of randomly chosen values from the set $\{-3,-1,1,3\}$. The signal $d(n)$ is passed through a linear channel [^2] In [48] the coefficient is written as 0.09 1; however, both the original [49] and the ‘other formulation’ [14] use 0.09, indicating that [48] may have a typographical error.:
$q(n)$ $=0.08d(n+2)-0.12d(n+1)+{}d(n)+0.18d(n-1)-0.1d(n-2)$ (6)
${}+0.09d(n-3)-0.05d(n-4)+0.04d(n-5)+{}0.03d(n-6)+0.01d(n-7)$
@@
where $v(n)$ is Gaussian white noise, applied to ensure that the signal-to-noise ratio of the output is between 12 and 32 decibels [25].333[25] specifies 12 to 32 dB; the supplementary material [48] specifies 16 to 32 dB. Most authors who use the benchmark [42, 40, 50, 51] use the lower bound of 12 dB. There are some other formulations; for example [14] has no noise term and adds an offset of +30. An example of such a signal, as well as the signal with noisy applied, can be found in figure 5.
Figure 5: A channel signal (bottom, blue), and the signal with nonlinear noise applied (top, orange).
One can also find certain tasks in the literature that are called ‘Channel Equalisation’, but do not follow the same methodology as the best known version of the benchmark. For example, Boccato et al. [52] use a similar task, but the signal is composed solely of the values –1, +1, and the evaluation is done via Average Mean Squared Error. Boccato et al. [53] follow this up by adding an unsupervised channel equalisation task, where the noiseless signal is not known. While this task has a similar purpose to the classic, supervised version of the task, the lack of “clean” signal used in training signifies that it is not an imitation task. This provides us with an interesting distinction between the form of the task (imitation of an unknown dynamical system, or not) and its function (channel equalisation, in both cases).
@@ 4.2.1 Unsupervised Channel Equalisation
An unsupervised version of the Channel Equalisation task exists [53]. The argument is that most distorted signals do not have a “clean” version that can be used to train the nonlinear filter, therefore it is useful to have a version of the task that uses unsupervised learning as opposed to imitation-based learning. Instead of training the received signal against the perfect, noiseless signal, the reservoir was trained against the information conveyed. The authors argue that, since the explicit signal itself was not used in training, this qualifies the task as an unsupervised learning task. As the ‘perfect signal’ is still known, however, the task is evaluated against that perfect signal.
@@ 4.3 Two-jointed arm: imitating unknown dynamics
The two-jointed arm task [54] involves controlling a two-jointed (robotic) arm in order to move it from point A to point B. Like the Channel Equalisation task, this task can be seen as an imitation of an unknown, perfect controller, of which only the outputs are known.
Joshi and Maass [55, 56] use this task in the context of Reservoir Computing. The training data used are points along the trajectory that the robotic arm was intended to follow. The robotic arm is then tested in a ‘closed loop’, where no training data is available, what we in section 3.1.1 refer to as ‘in-situ’ testing.
@@
where $\boldsymbol{\theta}=[\theta_{1},\theta_{2}]^{T}$ are the angles of the arm’s joints, $\boldsymbol{\tau}=[\tau_{1},\tau_{2}]^{T}$ are the joint input torques, $\mathbf{H}$ is a $2\times 2$ inertia matrix, and $\mathbf{C}$ is a $2\times 2$ matrix of the Coriolis and centripetal terms. The reservoir is then given a target trajectory, for which it must imitate a ‘perfect’ controller in a model-free manner. See [55, 56] for details.
@@ 4.4 Van der Pol oscillator: imitating unknown dynamics of a controller
The Van der Pol (VdP) oscillator was first described by the Dutch physicist Balthazar Van der Pol in the 1920s [57] and later employed to model the oscillations of the heart [58]. It is particularly useful for modelling systems that exhibit non-linear behaviour, such as relaxation oscillations and self-sustained oscillations. Within the context of machine learning, the VdP oscillator is commonly used to test the effectiveness of neural networks [59, 60]. The VdP oscillator can be applied as the substrate of physical reservoir computing [61].
@@
where $x$ is the state variable, $t$ is time, and $\mu$ is a scalar parameter that dictates the nonlinearity and the strength of the damping. $F(t)$ is the external force applied to the VdP system. In the undriven VdP system the external force term is zero $F(t)=0$, with the phase plane trajectory as shown in Figure 6.
@@
In a recent study, the VdP oscillator is used as a model-free control task where a reservoir is trained to produce an external force to make the VdP system produce a circle [23, 62].
@@ 4.5 Pole balancing : unsupervised learning
In the field of Reservoir Computing, this task is primarily used to evaluate unsupervised learning in reservoirs. It has been used for exploring different methods of training reservoirs, such as neuroevolution [63, 64] and NEAT [65]. One could also implement this task as an imitation task, and have the reservoir imitate the unknown dynamics of a perfect controller.
@@ 5 Dynamics prediction tasks
In this section, we describe the systems commonly used as prediction benchmarks in Reservoir Computing. A greater variety of datasets that can be used for benchmark tasks can be found in Gilpin [24]’s database of chaotic dynamical systems, all specified according to guidelines proposed by Gebru et al. [66].
@@ 5.1 Mackey–Glass equation : known dynamics
The Mackey–Glass benchmark is derived from a delay differential equation introduced to model certain physiological systems [67]. [^5] This delay differential equation is also used elsewhere in Reservoir Computing to create a nonlinear node in delay line reservoirs (see, for example, [13]). This other usage is not examined here. These models characterise illness through changes to the variables that lead to these systems turning chaotic. The authors examine various models, one for the breathing patterns of patients with Cheyne-Stokes illnesses, and two variants for the growth rates of white blood cells in leukaemia patients.
The latter model is used as the basis for the benchmark; [67, eqn.4b] states:
@@
Normalising the concentration with respect to the baseline, $x_{t}=P_{t}/\theta$, gives the more usual form of the Mackey–Glass chaotic equation [68, eqn.1]:
@@
Here $x$ is a dimensionless normalised quantity, expressed in ‘units’ of the parameter $\theta$. $\theta$ is still a parameter of the model, but is now implicit in the equation: it provides the scale for the dependent variable $x$. Eqn.11 has chaotic dynamics for $\beta=2$, $n=9.65$, $\gamma=1$, $\tau=2$ with an initial condition of $x=0.5$ for $t<0$ [68]. [^6] The text of [68] states $\beta=0.2,\gamma=0.1,\tau=2$, but the captions of its figures 1 and 2 state $\beta=2,\gamma=1,\tau=2$. Figure 7 here favours the values from the captions. Equivalently, one could use $\beta=0.2,\gamma=0.1,\tau=20$.
@@
Top: $[\beta,n,\gamma,\tau]$ values from [68]:
$\bullet~{}[0.2,9.65,0.1,2]$ : transient to a point attractor
$\bullet~{}[2,8.5,1,2]$ : periodic behaviour [their fig.5]
$\bullet~{}[2,9.65,1,2]$ : chaotic behaviour [their fig.2,fig.7]
Bottom: $[\beta,n,\gamma,\tau]$ values from [1]:
$\bullet~{}[0.2,10,0.1,16]$ : periodic behaviour
$\bullet~{}[0.2,10,0.1,17]$ : mildly chaotic behaviour
$\bullet~{}[0.2,10,0.1,30]$ : wildly chaotic behaviour
@@
Jaeger [1] states that the Mackey–Glass system is often used for learning dynamical systems from data, ‘invariably’ with parameter values of $\beta=0.2,n=10,\gamma=0.1$, and that with these values the system is chaotic for $\tau>16.8$; see [69] for a detailed analysis. Jaeger [1] investigates $\tau=17$ (mildly chaotic) and $\tau=30$ (which he dubs ‘wildly’ chaotic). See figure 7.
The parameter values used in [68] and [1] appear to differ considerably (figure 7); however, timescale rescaling shows they are comparable. There are three timescales in eqn.11: $1/\beta$ (production, or growth, timescale); $1/\gamma$ (decay timescale), and $\tau$ (time delay). If we normalise these with respect to the decay timescale, we have two independent timescales: $\gamma/\beta$ and $\gamma\tau$. [68] uses $\gamma/\beta=1/2$, $\gamma\tau=2$; [1] uses $\gamma/\beta=1/2$, $\gamma\tau=1.7,3.0$. [68] explores the transition between periodic and chaotic behaviour using fixed timescales and varying the non-linearity parameter $n$ (the timescales have biological meaning, whereas the non-linearity is a parameter of the model). [1] fixes the non-linearity at $n=10$ and instead varies the delay feedback timescale $\tau$; here the model is being used in a more abstract manner, and varying the time delay is a standard approach to investigating chaotic behaviours in DDEs [70].
Eqn.11 describes a continuous system. Jaeger [1] describes a discretisation process for converting it to a discrete system suitable for generating time series data.
  • Discretise eqn.11 using $dx/dt\approx(x(t+\Delta t)-x(t))/\Delta t$:
@@
  • Take $1/\Delta t=N$, an integer ([1] uses $N=10$), to produce the time series $\hat{x}(t),\hat{x}(t+\Delta t),\hat{x}(t+2\Delta t),\ldots\hat{x}(t+(N-1)\Delta t),\hat{x}(t+1),\ldots$
@@
Using this discretisation method, [1] generates two sequences for each value of $\tau=17$ and $30$, one sequence of length 3,000 and of 21,000 (four sequences in total). These form the combined washout and training sequences. The ESN is trained on this data, then tested on its ability to predict the 84th output value, compared to the expected 84th value, computing the relevant Normalised Mean Squared Error (NMSRE84) of the value over multiple runs.
Mackey–Glass is used as a benchmark by Jaeger again in [25], this time with a single 3,000-step training sequence, and the NMSRE of the 84th subsequent value taken over 100 runs. The $\tau$ values used are not specified. Holzmann and Hauser [39] also use the system as a benchmark, using a method almost identical to [1], though with the NMSRE taken over 100 runs, and an additional NMSRE of the 120th value used for the training sequence of length 21,000. Roeschies and Igel [71] base their method on [1], this time using a single sequence of length 3,000 with $\tau=17$. One other notable difference is the use of a washout of 100 values, whereas prior works have used washouts of 1,000. [41] and [72] depart from this traditional setup in order to propose their own. These various different uses are detailed in table 3.
@@
[71] 1 3000 17 100 84 50
[41] 1 2000 ? ? 2000 10
[72] 1 500 18 ? 50–52 ?
Table 3: A summary of different parameters used in the literature for the Mackey–Glass prediction benchmark. Values marked with ‘?’ are not reported in the cited work. [41] uses NMSE rather than NMSRE.